solve the equation:
81^3x=27^x+2

Answer by Alan3354(69443)

81^3x=27^x+2
Assuming x+2 is the exponent:
---------
81 = 3^4
27 = 3^3
-----------
3^12x = 3^(3x+6)
12x = 3x+6
9x = 6
x = 2/3